OpenStudy (gabylovesyou):

10. Find the 9th partial sum of the summation of 3i minus 4, from i equals 1 to infinity. 102 99 205 200

OpenStudy (gabylovesyou):

@mathmale @jigglypuff314

OpenStudy (anonymous):

For the sum\[\sum_{i=1}^{n}3i-4\] the partial sum is as follows \[\sum_{i=1}^{n}3i-4=\frac{ 3 }{ 2 }n^2-\frac{ 5 }{ 2 }n\] So the 9th partial sum is just n=9 \[\frac{ 3 }{ 2 }9^2-\frac{ 5 }{ 2 }9=99\]